4(1-3x)-2(x-4)=2(-8x+5)
Answer by Boreal(15235)

4(1-3x)-2(x-4)=2(-8x+5)
4-12x-2x+8=-16x+10. Distribute and watch the signs
12-14x=-16x+10 add 16 x to both sides
12+2x=10 subtract 12 from both sides
2x=-2 divide both sides by 2
x=-1
4(1+3)-2(-5)=2(13)
16+10=26